# Deep Multi-Task Learning Has Low Amortized Intrinsic Dimensionality

### v1

@@ Abstract
Deep learning methods are known to generalize well from training to future data, even in an overparametrized regime, where they could easily overfit. One explanation for this phenomenon is that even when their ambient dimensionality, (i.e. the number of parameters) is large, the models’ intrinsic dimensionality is small, i.e. their learning takes place in a small subspace of all possible weight configurations.
@@ 1 Introduction
Only recently has machine learning theory started to catch up with practical developments. Instead of overly pessimistic generalization bounds based on traditional complexity measures, such as VC dimension (Vapnik & Chervonenkis, 1971), Rademacher complexity (Bartlett & Mendelson, 2002) or overly pessimistic PAC-Bayesian estimates (McAllester, 1998; Alquier, 2024), non-vacuous bounds for neural networks were derived based on compact encodings. At their core lies the insight that—despite their usually highly overparametrized form—the space of functions learned by deep models has a rather low intrinsic dimensionality when they are trained on real-world data (Li et al., 2018). As a consequence, deep models are highly compressible, in the sense that a much smaller number of values (or even bits) suffices to describe the learned function compared to what one would obtain by simply counting the number of their parameters (Zhou et al., 2019; Lotfi et al., 2022).
In this work, we build on these insights to tackle another phenomenon so far not explained by generalization theory: the remarkable ability of deep networks for multi-task learning. In a setting where multiple related tasks are meant to be learned, high-accuracy deep models that generalize well can be trained from even less training data per task than in the standard single-task setting. Our contributions are the following:
  • We introduce a new way to express and quantify the low intrinsic dimensionality of deep multi-task learning models. Specifically, high-dimensional models are parametrized as linear combinations of a small number of basis elements, which themselves are trained in a low-dimensional random subspace of the original network parametrization.
  • We then demonstrate experimentally that in deep multi-task learning a much smaller effective dimension per task can suffice compared to single-task learning. Consequently, when learning related tasks, high-accuracy models can be characterized by a very small number of per-task parameters, which suggests an explanation for the strong generalization ability of deep MTL in this setting.
@@ 2 Background
In a multi-task setting, several learning tasks, $t_{1},\dots,t_{n}$, are meant to be solved. A multi-task learning (MTL) algorithm has access to all training sets, $S_{1},\dots,S_{n}$ and it is meant to output one model per task, $f_{1},\dots,f_{n}$. The hope is that if the tasks are related to each other in some form, a smaller overall risk $\mathcal{R}(f_{1},\dots,f_{n})=\frac{1}{n}\sum_{i=1}^{n}\mathcal{R}(f_{i})$ can be achieved than if learning each task separately (Caruana, 1997; Thrun & Pratt, 1998; Baxter, 2000).
@@ 2.1 Intrinsic Dimensionality of Neural Networks
where $\theta_{0}$ is a random initialization that helps optimization (Glorot & Bengio, 2010), and $w\in\mathbb{R}^{d}$ is a learnable low-dimensional vector, which is expanded to full dimensionality by multiplication with a fixed large matrix, $P\in\mathbb{R}^{D\times d}$, with uniformly sampled random entries. The authors found that $d$ could usually be chosen an order of magnitude smaller than $D$ without sacrificing model accuracy. Later, Aghajanyan et al. (2021) made similar similar observations in the context of model fine-tuning, and Lotfi and co-authors used this effective dimension to establish non-vacuous generalization bounds for learning in the setting of single-task prediction (Lotfi et al., 2022) and generation (Lotfi et al., 2024).
@@ 3 Amortized Intrinsic Dimensionality for Deep Multi-Task Learning
Our main claim in this work is that the success of deep multitask learning can be explained by their ability to effectively exploit and encode the relatedness between tasks in the form of a low intrinsic multi-task representation dimension.
To quantify this statement, we extend the definition of intrinsic dimensionality from a single-task setting, where no sharing of information between models takes place, to the multi-task setting, where sharing information between models can occur. Specifically, we introduce the amortized intrinsic dimension (AID) that extends the parametrization (1) in a hierarchical way. Instead of learning in a fully random subspace, we divide the multitask learning task into two components: learning the subspace (for which data of all tasks can be exploited), and learning per-task models within the subspace, for which only the data of the respective task is relevant. Both processes happen simultaneously in an end-to-end fashion, thereby making full use of the high-dimensional ambient space during the optimization progress.
@@
Table 1: Intrinsic dimensions for single-task learning ($\text{ID}_{90}$) and multi-task learning ($\text{AID}_{90}$) at fixed target accuracies $\text{acc}_{90}$ for different datasets ($n$: number of tasks, $m$: average number of examples per task, $L$: number of classes, $l$, $k$: subspace dimension and number of basis vectors in representation (2)) resulting in the reported $\text{AID}_{90}$.
@@
$n$ / $m$ / $L$ 30 / 2000 / 10 30 / 2000 / 10 60 / 900 / 2 60 / 2000 / 2 100 / 453 / 3 100 / 450 / 10
target accuracy $\text{acc}_{90}$ 85% 87% 65% 75% 63% 40%
# model parameters 21840 21840 11810 13730 121182 128832
single-task $\text{ID}_{90}$ 400 300 50 50 200 1500
multi-task $\text{AID}_{90}$ 31.6 166.6 10 10 12 36
($l$, $k$) (65, 10) (70, 50) (60, 5) (60, 5) (20, 10) (80, 20)
Overall, this formulation uses $l\cdot k$ parameters to learn an expansion matrix that can capture information that is shared across all tasks. In addition, for each of the $n$ tasks, $k$ additional parameters are used to describe a good model within the shared subspace. Consequently, the total number of training parameters which we call multi-task intrinsic dimensionality is $lk+nk$. Expressing this on the per-task level, we obtain the following definition.
Definition 3.1.
In the setting described above, we define the multi-task amortized intrinsic dimensionality (AID) as $\frac{lk}{n}+k$.
In comparison, learning each task separately using the representation (1) would require $d$ parameter per task. Based on these considerations, we formulate our central scientific hypothesis:
Hypothesis.
Deep MTL can find models of high accuracy with much smaller amortized intrinsic dimensionality than the average intrinsic dimensionality required for learning the same tasks separately: $\frac{lk}{n}+k\ll d$.
@@ 4 Determining the AID for Real-World Tasks
Figure 1: Amortized intrinsic dimensionality for multi-task learning on the MNIST SP dataset ($n=10,20,\dots,100$ tasks, $m=600$ samples per task). The more tasks are available, the smaller the number of parameters per task required to achieve the same accuracy.
In this section, we provide evidence for our hypothesis by numerically estimating AID values across different settings and comparing them with their counterparts for single-task learning.
Following previous works (Li et al., 2018) the measure of comparison is the smallest dimension required to reach 90% of the average accuracy achieved by training a fully parametrized (i.e. not in the representation (1)) single-task models of the same architecture. We call this accuracy value $\text{acc}_{90}$ and the dimension required to reach it $\text{ID}_{90}$ in the single-task and $\text{AID}_{90}$ in the multi-task setting. To experimentally estimate the intrinsic dimensions, we rely on six standard multi-task learning benchmarks: MNIST SP, MNIST PL, split-CIFAR10 and split-CIFAR100 are multi-class image datasets. We use networks with convolutional architecture for them. Products, which consists of vectorial embeddings of text data, and Folktables, which contains tabular data, are binary classification tasks, for which we use fully connected architectures. For details of the dataset and network architectures, see Table 1 and Section 7.
Table 1 reports the intrinsic dimensions that result from our experiments. It clearly supports our hypothesis: in all tested cases, the amortized intrinsic dimensionality in the multi-task setting is smaller than the per-task intrinsic dimensionality when training tasks individually, sometimes dramatically so. For example, to train single-task models on the 30 MNIST SP tasks, it suffices to work in a 400-dimensional subspace of the models’ overall 21840-dimensional parameter space. For multi-task learning in representation (3), it even suffices to learn a basis of dimension $k=10$, where each element is learned within a subspace of dimension $l=65$ of the model parameters. Consequently, MTL represents all 30 models by just $10\cdot 65+30\cdot 10=950$ parameters, resulting in an amortized intrinsic dimension of just $950/30\approx 32$. The other datasets show similar trends: for the Folktables and Products datasets, the intrinsic dimension is reduced from 50 to 10 in both cases. For split-CIFAR10 and split-CIFAR100, the intrinsic dimensions drop from 200 and 1500 to 12 and 36, respectively. The smallest gains we observe for MNIST PL, where the amortized intrinsic dimensions are slightly more than half of the intrinsic dimension of single-task learning.
@@
5 Generalization Guarantees
The result in the previous section suggests that using the representation (2), even quite complex models can be represented with only a quite small number of values. This suggests that the good generalization properties of deep multi-task learning might be explainable by generalization bounds that exploit the representation (3).
@@ 5.1 Background and Related Work
Before we formulate our results, we remind the reader of some core concepts and results from the literature.
Definition 5.1.
@@
Definition 5.2.
For any model set, $\mathcal{F}$, an encoder with a base set $\mathcal{Z}=\mathcal{F}$ is called a model encoder.
@@
Theorem 5.3 (Shalev-Shwartz & Ben-David (2014), Theorem 7.7).
@@
Theorem 5.3 states that models can be expected to generalize well, i.e., have a smaller difference between their training risk and expected risk, if their encodings are short. Specifically, the dominant term in the complexity term of (4) is $\sqrt{\frac{l_{E}(f)\log 2}{2m}}$. Consequently, for the bound to become non-vacuous (i.e. the right-hand side to be less than 1), in particular, the encoding length must not be much larger than the number of training examples.
5.2 New Guarantees for Multi-Task Learning
We now introduce a generalization of Theorem 5.3 to the multi-task situation. First, we define two types of encoders.
Definition 5.4.
For any model set, $\mathcal{F}$, an encoder with base set $\mathcal{Z}=\bigcup_{n=1}^{\infty}\mathcal{F}^{n}$ (e.g. arbitrary length tuples of models) we call a multi-task model encoder.
Multi-task model encoders generalize (single-task) model encoders in the sense of Definition 5.2 by allowing more than one model to be encoded at the same time. By exploiting redundancies between the models, shorter encoding lengths can be achievable. For example, if multiple models share certain parts, such as feature extraction layers, those would only have to be encoded once.
Any such construction will work better (i.e. results shorter length functions) in some situations, but potentially worse in another situation, depending on if and how the tasks to be learned are related. For multi-task learning, it would therefore be beneficial if one could choose the encoding scheme in a data-dependent way. To formalize this, we introduce meta-encoders.
Definition 5.5.
An encoder is called a meta-encoder if its base set is a set of multi-task model encoders.
While meta-encoders are less intuitive objects than, e.g., standard model encoders, they nevertheless serve an important purpose in our analysis, because they allow formalizing the concept of shared versus task-specific information. For example, if all models reside in a low-dimensional subspace of the parameter space, the meta-encoder could encode the subspace, while the multi-task model encoder only encodes the representation of the models within that subspace.
@@ 5.2 New Guarantees for Multi-Task Learning
Theorem 5.6.
Let $\mathcal{E}$ be a set of multi-task model encodings with length functions $l_{E}:\bigcup_{n=1}^{\infty}\mathcal{F}^{n}\rightarrow\mathbb{N}$ for $E\in\mathcal{E}$. For any meta-encoding with length function $l:\mathcal{E}\rightarrow\mathbb{N}$, and any $\delta>0$, the following inequality holds with probability at least $1-\delta$ over the sampling of the training data: for all $E\in\mathcal{E}$ and for all $f_{1},...,f_{n}\in\mathcal{F}$
$\mathcal{R}(f_{1},...,f_{n})\leq\widehat{\mathcal{R}}(f_{1},...,f_{n})$ (5)
$\quad+\sqrt{\frac{(l(E)+l_{E}(f_{1},...,f_{n}))\log(2)+\log\frac{1}{\delta}}{2mn}}.$
Theorem 5.6 directly generalizes Theorem 5.3 to the multi-task situation.
@@
theoremfastrate In the settings of Theorem 5.6, for any multi-task model encoding, any meta-encoding, and any $\delta>0$, the following holds with probability at least $1-\delta$ over the sampling of the training data:
$\operatorname{\mathbf{kl}}\Big{(}\widehat{\mathcal{R}}(f_{1},\dots,f_{n})|\mathcal{R}(f_{1},\dots,f_{n})\Big{)}$
$\leq\frac{(l(E)+l_{E}(f_{1},\dots,f_{n}))\log(2)+\log(\frac{2\sqrt{mn}}{\delta})}{mn},$ (6)
where
$\operatorname{\mathbf{kl}}(q|p)=q\log\frac{q}{p}+(1-q)\log\frac{1-q}{1-p}$ (7)
is the Kullback–Leibler divergence between Bernoulli distributions with mean $q$ and $p$. Fast-rate bounds tend to be harder to prove and interpret, but they offer tighter results when the empirical risk is small. Specifically, the complexity term in Theorem 5.6 scales like $O(\frac{1}{mn})$ when the empirical risk is zero, whereas in Theorem 5.6 the scaling behaviour is always $O(\frac{1}{\sqrt{mn}})$. For a more detailed discussion, see Appendix A.

### v2

# From Low Intrinsic Dimensionality to Non-Vacuous Generalization Bounds in Deep Multi-Task Learning

@@ Abstract
Deep learning methods are known to generalize well from training to future data, even in an overparametrized regime, where they could easily overfit. One explanation for this phenomenon is that even when their ambient dimensionality, (i.e. the number of parameters) is large, the models’ intrinsic dimensionality is small; specifically, their learning takes place in a small subspace of all possible weight configurations.
@@ 1 Introduction
Only recently has machine learning theory started to catch up with practical developments. Instead of overly pessimistic generalization bounds based on traditional complexity measures, such as VC dimension (Vapnik & Chervonenkis, 1971), Rademacher complexity (Bartlett & Mendelson, 2002) or often loose PAC-Bayesian estimates (McAllester, 1998; Alquier, 2024), non-vacuous bounds for neural networks were derived based on compact encodings (Zhou et al., 2019; Lotfi et al., 2022). At their core lies the insight that—despite their usually highly overparametrized form—the space of functions learned by deep models has a rather low intrinsic dimensionality when they are trained on real-world data (Li et al., 2018). As a consequence, deep models are highly compressible, in the sense that a much smaller number of values (or even bits) suffices to describe the learned function compared to what one would obtain by simply counting the number of their parameters.
In this work, we leverage these insights to tackle another phenomenon currently lacking satisfactory generalization theory: the remarkable ability of deep networks for multi-task learning. In a setting where multiple related tasks are meant to be learned, high-accuracy deep models that generalize well can be trained from even less training data per task than in the standard single-task setting. Our contributions are the following:
  • We introduce a new way to parametrize and analyze a class of deep multi-task learning models that allows us to quantify their intrinsic dimensionality. Specifically, high-dimensional models are learned as linear combinations of a small number of basis elements, which themselves are trained in a low-dimensional random subspace of the original network parametrization.
  • We then demonstrate experimentally that for such networks a much smaller effective dimension per task can suffice compared to single-task learning. Consequently, when learning related tasks, high-accuracy models can be characterized by a very small number of per-task parameters, which suggests an explanation for the strong generalization ability of deep MTL in this setting.
@@ 2 Background
In a multi-task setting, several learning tasks, $t_{1},\dots,t_{n}$, are meant to be solved. A multi-task learning (MTL) algorithm has access to all training sets, $S_{1},\dots,S_{n}$ and it is meant to output one model per task, $f_{1},\dots,f_{n}$ [^1] In practice, one might enforce that these models share certain components, e.g. a token embedding layer or a feature extraction stage. In this work, we do not make any such a prior assumptions and let the multi-task learning algorithm decide in what form to share parameters, if any.. The expectation is that if the tasks are related to each other in some form, a smaller overall risk $\mathcal{R}(f_{1},\dots,f_{n})=\frac{1}{n}\sum_{i=1}^{n}\mathcal{R}(f_{i})$ can be achieved than if each task is learned separately (Caruana, 1997; Thrun & Pratt, 1998; Baxter, 2000).
@@ 2.1 Intrinsic dimensionality of neural networks
where $\theta_{0}$ is a random initialization that helps optimization (Glorot & Bengio, 2010), and $w\in\mathbb{R}^{d}$ is a learnable low-dimensional vector, which is expanded to full dimensionality by multiplication with a fixed large matrix, $P\in\mathbb{R}^{D\times d}$, with uniformly sampled random entries. The authors introduced a procedure in which $d$ is determined such that a certain accuracy level is achieved, typically $90$% of an ordinarily-trained full-rank model. As such, this effective dimension is a property of the model as well as the data, and the authors found it could usually be chosen orders of magnitude smaller than the ambient dimension (number of model parameters) $D$. Later, Aghajanyan et al. (2021) made similar observations in the context of model fine-tuning. Lotfi and co-authors used the effective dimension to establish non-vacuous generalization bounds for learning in the setting of single-task prediction (Lotfi et al., 2022) and generative modeling (Lotfi et al., 2024).
@@ 3 Amortized intrinsic dimensionality for deep multi-task learning
Our main claim in this work is that the success of deep multi-task learning can be explained by its ability to effectively express and exploit the relatedness between tasks in a low-dimensional subspace of the parametrization space.
To quantify this statement, we extend the definition of intrinsic dimensionality from a single-task setting, where no sharing of information between models takes place, to the multi-task setting, where sharing information between models can occur. Specifically, we introduce the amortized intrinsic dimension (AID) that extends the parametrization (1) in a hierarchical way. Instead of learning in a fully random subspace, we divide the multi-task learning task into two components: learning the subspace (for which data of all tasks can be exploited), and learning per-task models within the subspace, for which only the data of the respective task is relevant. Both processes occur simultaneously in an end-to-end fashion, thereby making full use of the high-dimensional model parameter space during the optimization progress.
@@
Overall, this formulation uses $l\cdot k$ parameters to learn an expansion matrix that can capture information that is shared across all tasks. In addition, for each of the $n$ tasks, $k$ additional parameters are used to describe a good model within the shared subspace. Consequently, the total number of training parameters in such a parametrization is $lk+nk$, i.e. $\frac{lk}{n}+k$ per task.
The following definition introduces the notions of intrinsic dimension and amortized intrinsic dimension based on the above construction.
Definition 1.
For a given model architecture and set of tasks, $t_{1},\dots,t_{n}$, assume that individual (single-task) models, $f_{i}$, are trained on each task. Let $A=\frac{1}{n}\sum_{i=1}^{n}\text{acc}(f_{i})$ be the average (validation or test) accuracy that these models achieve on their respective tasks. We define
  • the (single-task) intrinsic dimensionality, $\text{ID}_{p}$, as the smallest value for $d$ in the $d$-dimensional expansion (1), such that training the corresponding low-rank models individually on each task results in an average across-tasks accuracy of at least $p$of $A$.
  • the (multi-task) amortized intrinsic dimensionality, $\text{AID}_{p}$, as the smallest value for $\frac{lk}{n}+k$ in an $(l,k)$-dimensional expansion (2)/(3), such that multi-task training the corresponding low-rank models results in an average across-tasks accuracy of at least $p$of $A$.
Following Li et al. (2018), in this work we always use the accuracy factor $p=90$, and we write ID and AID as shorthand notations for $\text{ID}_{90}$ and and $\text{AID}_{90}$, respectively.
We now formulate our central scientific hypothesis:
Hypothesis.
Deep MTL can find models of high accuracy with much smaller amortized intrinsic dimensionality than the intrinsic dimensionality required for learning the same tasks separately: $\text{AID}_{90}\ll\text{ID}_{90}$
Note that the validity of this hypothesis is not obvious, and its correctness depends on the model architecture as well as set of tasks to be learned. For example, if the tasks are completely unrelated such that no shared subspace is expressive enough to learn good models for all of them, one likely would need one basis element per tasks ($k=n$), and the basis would need the same dimensionality as for single-task learning ($l=\text{ID}_{90}$). Consequently, $\text{AID}_{90}=\text{ID}_{90}+n$, i.e. the multi-task parametrization is not more parameter-efficient than the single-task one. In the opposite extreme, if all tasks are identical, a single $\text{ID}_{90}$-dimensional model suffices to represent all solutions, i.e. $k=1$ and $l=\text{ID}_{90}$, such that $\text{AID}_{90}=\frac{\text{ID}_{90}}{n}+1$, i.e. the amortized intrinsic dimensionality of multi-task learning shrinks quickly with the number of available tasks. For real-world settings, where tasks are related but not identical, we expect that $k$ and $l$ will have to grow with the number of tasks, but at a sublinear speed compared to $n$.
Table 1: Intrinsic dimensions for single-task learning ($\text{ID}_{90}$) and multi-task learning ($\text{AID}_{90}$) at fixed target accuracies $\text{acc}_{90}$ for different datasets and model architectures (with $D$ parameters) resulting in the reported $\text{ID}_{90}$ and $\text{AID}_{90}$.
@@
model ConvNet ConvNet MLP MLP ConvNet ViT ConvNet ViT
$n$ / $m$ 30 / 2000 30 / 2000 60 / 900 60 / 2000 100 / 453 30 / 1248 100 / 450 30 / 1250
$D$ 21840 21840 11810 13730 121182 5526346 128832 5543716
$\text{acc}_{90}$ 85% 87% 65% 75% 63% 80% 40% 65%
$\text{ID}_{90}$ 400 300 50 50 200 200 1500 550
$\text{AID}_{90}$ 31.6 166.6 10 10 12 26.7 36 100
($l$, $k$) (65, 10) (70, 50) (60, 5) (60, 5) (20, 10) (50, 10) (80, 20) (70, 30)
@@ 4 Determining the AID for real-world tasks
In this section, we provide evidence for our hypothesis by numerically estimating AID and ID values across different settings. We rely on six standard multi-task learning benchmarks: MNIST Shuffled Pixels (SP), MNIST Permuted Labels (PL), split-CIFAR10 and split-CIFAR100 are multi-class image datasets, and we report results for ConvNets and Vision Transformers for them. Products, which consists of vectorial embeddings of text data, and Folktables, which contains tabular data, are binary classification tasks, for which we use fully connected architectures. For more details on the datasets and network architectures, see Appendix B.
Table 1 reports the intrinsic dimensions obtained in our experiments. It clearly supports our hypothesis: in all tested cases, the amortized intrinsic dimensionality in the multi-task setting is smaller than the per-task intrinsic dimensionality when training tasks individually, sometimes dramatically so. For example, to train single-task models on the 30 MNIST SP tasks, it suffices to work in a 400-dimensional subspace of the models’ overall 21840-dimensional parameter space. For multi-task learning in representation (3), it even suffices to learn a basis of dimension $k=10$, where each element is learned within a subspace of dimension $l=65$ of the model parameters. Consequently, MTL represents all 30 models by just $10\cdot 65+30\cdot 10=950$ parameters, resulting in an amortized intrinsic dimension of just $950/30\approx 32$. The other datasets show similar trends: for the Folktables and Products datasets, the intrinsic dimension is reduced from 50 to 10 in both cases. For split-CIFAR10 and split-CIFAR100, the intrinsic dimensions drop from 200 and 1500 to 12 and 36, respectively. The smallest gain we observe is for MNIST PL, where the amortized intrinsic dimensions are slightly more than half of the intrinsic dimension of single-task learning.
@@
Note that the provided results in Table 1, show the amortized intrinsic dimensionality for our model design. Therefore, the reported values can be seen as upper bounds on the minimum possible amortized intrinsic dimension of multi-task learning. Choosing a specific multi-task design based on prior knowledge about the relatedness of the tasks could push the numbers in Table 1 even lower. However, even without any such additional information, our reported results clearly confirm our hypothesis by showing a substantial reduction in $\text{AID}_{90}$ compared to $\text{ID}_{90}$.
Figure 1: Amortized intrinsic dimensionality, $\text{AID}_{90}$, for multi-task learning on the MNIST PL (left) and MNIST SP (right) datasets ($n=10,20,\dots,100$ tasks, $m=600$ samples per task). The more tasks are available, the fewer learnable parameters per task are required.
5 Non-vacuous generalization guarantees for multi-task learning
The result in the previous section suggests that using the representation (3), even quite complex models can be represented with only a quite small number of values. This suggests that the good generalization properties of deep multi-task learning might be explainable by generalization bounds that exploit this fact.
@@ 5.1 Background and related work
Before formulating our results, we remind the reader of some classical concepts and results.
Definition 2.
@@
Definition 3.
For any set of models, $\mathcal{F}$, an encoder with a base set $\mathcal{Z}=\mathcal{F}$ is called a model encoder.
@@
Theorem 1 (Shalev-Shwartz & Ben-David (2014), Theorem 7.7).
@@
Theorem 1 states that models can be expected to generalize well, i.e., have a smaller difference between their training risk and expected risk, if their encodings are short. Specifically, the dominant term in the complexity term of (4) is $\sqrt{\frac{l_{E}(f)\log 2}{2m}}$. Consequently, for the bound to become non-vacuous (i.e. the right-hand side to be less than 1), in particular, the encoding length must not be much larger than the number of training examples.
5.2 Compression-based guarantees for multi-task learning
We now introduce a generalization of Theorem 1 to the multi-task situation.
First, we define two types of encoders, that allow us to formalize the concept of shared versus task-specific information in multi-task learning.
Definition 4.
For any model set, $\mathcal{F}$, we define two type of encoders:
  • meta-encoder: an encoder whose its base set is a set of global parameters $E\in\mathcal{E}$ representing shared information between tasks.
  • multi-task (model) encoder: an encoder with base set $\mathcal{Z}=\bigcup_{n=1}^{\infty}\mathcal{F}^{n}$ (e.g. arbitrary length tuples of models) that, given a global parameter $E$, encodes a tuple of models with the length $l_{E}(f_{1},\dots,f_{n})$ i.e. encodes the task-specific information.
These encoders generalize (single-task) model encoders in the sense of Definition 3 by allowing more than one model to be encoded at the same time. By exploiting redundancies between the models, shorter encoding lengths can be achievable. For example, if multiple models share certain parts, such as feature extraction layers, those would only have to be encoded once by meta-encoder while multi-task encoder encodes the remaining parts. Additionally, as we will explain in in Section 5.3.1 multi-task encoders can capture unstructured redundancies between task-specific parts to reduce the encoding lengths.
@@ 5.2 Compression-based guarantees for multi-task learning
{restatable}
theoremslowrate Let $\mathcal{E}$ be a set of global parameters, and let $l:\mathcal{E}\rightarrow\mathbb{N}$ be the length function of the meta-encoder. For each $E\in\mathcal{E}$, let $l_{E}:\bigcup_{n=1}^{\infty}\mathcal{F}^{n}\rightarrow\mathbb{N}$ be the length function of the multi-task encoder given $E$. For any $\delta>0$, with probability at least $1-\delta$ over the sampling of the training data for all $E\in\mathcal{E}$ and for all $f_{1},...,f_{n}\in\mathcal{F}$ the following inequality holds:
$\mathcal{R}(f_{1},...,f_{n})$$\leq\widehat{\mathcal{R}}(f_{1},...,f_{n})+\sqrt{\frac{(l(E)+l_{E}(f_{1},...,f_{n}))\log(2)+\log\frac{1}{\delta}}{2mn}}.$ (5)
Definition 4 directly generalizes Theorem 1 to the multi-task situation.
@@
theoremfastrate In the settings of Definition 4, for any multi-task model encoding, any meta-encoding, and any $\delta>0$, the following holds with probability at least $1-\delta$ over the sampling of the training data:
$\operatorname{\mathbf{kl}}\Big{(}\widehat{\mathcal{R}}(f_{1},\dots,f_{n})|\mathcal{R}(f_{1},\dots,f_{n})\Big{)}$ $\leq\frac{(l(E)+l_{E}(f_{1},\dots,f_{n}))\log(2)+\log(\frac{2\sqrt{mn}}{\delta})}{mn},$ (6)
where $\operatorname{\mathbf{kl}}(q|p)=q\log\frac{q}{p}+(1-q)\log\frac{1-q}{1-p}$ is the Kullback-Leibler divergence between Bernoulli distributions with mean $q$ and $p$. Fast-rate bounds tend to be harder to prove and interpret, but they offer tighter results when the empirical risk is small. Specifically, the complexity term in Definition 4 scales like $O(\frac{1}{mn})$ when the empirical risk is zero, whereas in Definition 4 the scaling behaviour is always $O(\frac{1}{\sqrt{mn}})$. For a more detailed discussion, see Appendix A.